8. Beatrice owns a cabinet installation business. Her accountant most likely includes which of the following costs on her financial statements?
A. wages Beatrice could earn building swing sets
B. the labor cost of Beatrice’s assistant cabinet installer.
C. dividends Beatrice's stock was earning before Beatrice sold her stock and used those funds to purchase a cabinet installation business.
D. All of the above
E. A and C, only.


9. A firm produces and sells potato chips. Last year it sold 3 million bags of potato chips at a price of $3 per bag. Total costs of production equal $5.5 million. If implicit costs are not equal to zero, then, for last year, the firm's
(x) total revenue was $9 million.
(y) economic profit was $3.5 million.
(z) accounting profit was more than $3.5 million
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (x) only


10. William decides to open his own business and earns $70,000 in accounting profit the first year. When deciding to open his own business, he turned down three separate job offers with annual salaries of $48,000, $52,000, and $55,000. What is William’s economic profit from running his own business?
A. –$55,000
B. $22,000
C. –$15,000
D. $15,000
E. $0
Answer #1

Part 8

Answer is B

The accountant will only include the explicit cost and not the implicit cost.

Part 9

Answer is E

Accounting Profit = Total Revenue - Explicit Cost

= P*Q = 3*3 = 9

= 9 million - 5.5 million

= 3.5 million

Since implicit cost are greater than 0, the economic profit should be less than 3.5 million.

Economic Profit = Accounting Profit - Implicit Cost  

Part 10

Answer is D

The implicit cost is the highest forgone salary

Economic Profit = Accounting Profit - Implicit Cost = 70000 - 55000 = 15000